# TianQuan-S2S: A Subseasonal-to-Seasonal Global Weather Model via Incorporate Climatology State

### v4

@@ 6 Conclusion
In this work, we present TianQuan-S2S, a novel framework that addresses the challenges in subseasonal-to-seasonal (S2S) forecasting, particularly the limitations of model collapse and the insufficient use of climate information. Our extensive experiments on the ERA5 dataset demonstrate that TianQuan-S2S outperforms both traditional numerical S2S systems and data-driven models, providing significant improvements in both deterministic and ensemble forecasting. Ablation studies have further substantiated the effectiveness of model designs and additional empirical analysis illustrates the superior performance across all lead times. In the future, we plan to ptimize our model framework to enable higher-resolution forecasts, and incorporate additional information such as land, ocean, and sea ice data to improve the learning capabilities of the model.
@@
Appendix A Notations
@@
Appendix B Model Details
@@ B.2 Hyperparameters
Appendix C Inference Details
@@
C.1 Forecast Strategy Research
@@
FuXi ✓ (ERA5) ✓ (Cascade) Each model 5d Up to 15d
ClimaX ✓ (CMIP6) ✓ (ClimaX-iter) 1, 3, 5, 7, 10, 14, 30d Up to 30d
@@
In addition, we chose 5-day intervals mainly to balance training cost and inference efficiency. Finer-grained models (e.g., daily intervals) significantly increase training time and computational resources.
To explore potential benefits, we also trained daily models for short lead times (within 15–20 days). The daily models take only the current timestep $t$ as input, while the 5-day models use a window of the past 5 timesteps (from t4t-4 to $t$) as input.
Table 8: Results for different intervals (1-day and 5-day).
Intervals 15 16 17 18 19 20
1-day 2.58 2.63 2.75 2.66 2.71 2.72
5-day 2.64 2.68 2.74 2.71 2.75 2.73
As shown in the Table 8, the daily models achieved slightly better accuracy than the 5-day models, but the improvement was marginal and did not bring large gains.
Hence, we opted for the 5-day interval setup as a more efficient and practical choice.
C.2 Layer to Lead Time Alignment in UD-ViT
To confirm that UD-ViT block indices map to specific lead times, we evaluate the 40-day model (8 blocks → 5 days per block). We unpatchify outputs after blocks 1, 2, 4, and 8, which correspond to lead times of 5, 10, 20, and 40 days, and compute the RMSE against the ground truth for T2m and Z500.
Table 9: RMSE values for different UD-ViT layers and lead times (lower is better).
Variables Layer ID 5 10 20 40
T2m Layer 1 1.68 2.63 2.94 3.241
Layer 2 2.24 2.45 2.82 3.089
Layer 4 2.58 2.68 2.73 3.125
Layer 8 2.61 2.71 2.841 2.998
Z500 Layer 1 358.6 830.5 884.3 924.5
Layer 2 560.2 782.1 874.6 915.4
Layer 4 760.2 826.4 869.2 910.9
Layer 8 852.3 861.2 876.2 904.5
The Table 9 verifies the alignment between UD-ViT block IDs and corresponding lead times, demonstrating that deeper layers correlate with longer-range forecasts.
Appendix D Employed Data
@@ D.4 Data Processing
Appendix E Quantitative evaluation
@@
Appendix F Additional Results
@@
As shown in the results 11, while these models perform reasonably well, our proposed TianQuan-S2S consistently achieves lower RMSE, particularly at longer lead times. This demonstrates the superiority of our architecture in S2S forecasting.

### v5

@@ 6 Conclusion and Future Work
In this work, we present TianQuan-S2S, a novel framework that addresses the challenges in subseasonal-to-seasonal (S2S) forecasting, particularly the limitations of model collapse and the insufficient use of climate information. Our extensive experiments on the ERA5 dataset demonstrate that TianQuan-S2S outperforms both traditional numerical S2S systems and data-driven models, providing significant improvements in both deterministic and ensemble forecasting. Ablation studies have further substantiated the effectiveness of model designs, and additional empirical analysis illustrates the superior performance across all lead times. In the future, we plan to optimize our model framework to enable higher-resolution forecasts and incorporate additional information such as land, ocean, and sea ice data to improve the learning capabilities of the model.
7 Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant T2125006; in part by GuangDong Basic and Applied Basic Research Foundation under Grant 2025A1515510016; in part by Shenzhen Science and Technology Program under Grant KCXFZ20240903093759004 and Grant KJZD20230923115106012, in part by the Open Research Fund of Pengcheng Laboratory under No.2025KF1B0040. Yang Liu is supported in part by the Postdoctoral Fellowship Scheme of The Chinese University of Hong Kong.
8 Reproducibility statement
We are committed to ensuring the reproducibility of our research. Source code and a README.md file with detailed instructions for data preparation and script execution are available at https://github.com/zhangminglang42/TianQuan and also provided in the Appendix. The appendix offers comprehensive details to support our claims. Appendix C describes the details and hyperparameters of our proposed TianQuan-S2S. In Appendix D, we further explored the differences in the prediction methods of each model and conducted extensive experiments on the prediction methods of our own model. Furthermore, Appendix E and F list in detail the sources and processing methods of the experimental data, and the calculation formulas of the indicators are also included.
@@ 8 Reproducibility statement
Appendix A The Use of Large Language Models (LLMs)
In this study, Large Language Models (LLMs) were used as an assistive tool to enhance the overall quality and clarity of the text. The primary research, analysis, and intellectual contributions remain solely the work of the authors. The key applications of LLMs in this study are as follows:
  • Text Refinement: LLMs were employed to enhance the grammatical accuracy, structure, and consistency of the manuscript. This involved refining sentence flow and ensuring consistent phrasing and tone across the entire paper.
  • Coherence and Structure: LLMs were used to organize and strengthen the logical flow of ideas. Suggestions were incorporated to improve transitions between sections and paragraphs, making the narrative more cohesive.
  • Idea Articulation: At various stages, LLMs acted as a sounding board, helping to articulate complex ideas more clearly and explore alternative expressions for the concepts already proposed by the authors.
All recommendations made by the LLM were critically reviewed, revised, and approved by the authors to ensure accuracy and alignment with the research intent. The final responsibility for the content lies solely with the authors.
Appendix B Notations
@@
Appendix C Model Details
@@ C.2 Hyperparameters
Appendix D Inference Details
@@
D.1 Forecast Strategy Research
@@
FuXi $\checkmark$ (ERA5) $\checkmark$ (Cascade) Each model 5d Up to 15d
ClimaX $\checkmark$ (CMIP6) $\checkmark$ (ClimaX-iter) 1, 3, 5, 7, 10, 14, 30d Up to 30d
@@
Appendix E Discussion
E.1 Consistency Analysis of Multi-Model Outputs
To address the concern regarding temporal discontinuities between adjacent lead times, our model primarily analyzes the continuity of the output at the two key points:
  • Residual connections for consistent representations: Residual connections in the Transformer preserve information across layers, helping the model maintain coherent representations from short to long leads and reducing abrupt shifts in the forecasts.
  • Smooth transitions and stability: Beyond climatology fusion, we add lead-time and absolute-time embeddings (days since 1 Jan 1970) so the model is explicitly aware of the target time, which improves temporal continuity and stability between adjacent lead-time forecasts.
Additionally, we further assess temporal continuity by computing the first-order differences between adjacent model outputs and analyzing mean/std/max in the Table 9.
Table 9: Mean, standard deviation, and maximum of first-order differences for different variables and lead-time ranges.
Metrics Vars 15–20 20–25 25–30 30–35 35–40 40–45
mean T2m 0.132 0.152 0.187 0.210 0.286 0.316
T850 0.199 0.206 0.227 0.239 0.302 0.329
Wind10 0.0138 0.0189 0.0196 0.0255 0.0353 0.0531
Z500 13.01 13.96 16.47 16.60 18.15 27.90
std T2m 1.973 1.986 2.432 2.596 2.603 2.716
T850 2.558 2.657 2.921 3.228 3.257 3.527
Wind10 0.971 1.128 1.180 1.377 1.414 1.497
Z500 176.2 173.7 203.5 182.3 204.5 226.2
max T2m 11.88 13.01 15.36 16.45 17.57 18.79
T850 17.43 22.09 27.16 28.22 28.75 28.82
Wind10 6.852 7.331 8.595 9.913 12.33 12.36
Z500 1031 1078 1252 1361 1322 1482
Note: K-(K+5) denotes the continuity metric between the K-day and (K+5)-day model predictions.
Overall, the mean, standard deviation, and maximum of the first-order differences increase with lead time across all variables, consistent with the gradual degradation of forecast skill at longer ranges. At the same time, the growth of these statistics across adjacent lead-time windows is smooth rather than abrupt, indicating that the fields evolve in a temporally continuous manner instead of exhibiting erratic, frame-to-frame jumps.
E.2 Challenge in Wind Forecasting
The challenge in accurately forecasting near-surface variables, such as Wind10, primarily stems from the extreme complexity of the underlying surface conditions and strong thermodynamic processes. Factors like heterogeneous terrain, land-sea contrast, and surface heating induce intense turbulence and high variability, making near-surface wind inherently less predictable than large-scale atmospheric variables.
We acknowledge that current data-driven models, including FuXi-S2S and TianQuan-S2S, are currently limited by the absence of high-resolution surface data and explicit solar radiation inputs, which are essential for capturing these fine-scale processes. This limitation explains why simple climatology, which smooths out subseasonal variability, can sometimes appear more stable in evaluation metrics for Wind10.
We fully agree on the importance of these inputs. The current model is evaluated primarily on lower-resolution data and does not yet incorporate other crucial Earth system components (e.g., land surface, ocean, and sea ice data), which significantly influence subseasonal climate variability.
We plan to extend the model to support higher-resolution inputs and integrate multi-source Earth observation data (including detailed surface and radiation information) to enhance its physical consistency and forecasting skill.
Appendix F Employed Data
@@ F.4 Data Processing
Appendix G Quantitative evaluation
@@
Appendix H Additional Results
@@
As shown in the Table 11, while these models perform reasonably well, our proposed TianQuan-S2S consistently achieves lower RMSE, particularly at longer lead times. This demonstrates the superiority of our architecture in S2S forecasting.
Table 12: T2m RMSE of ablation studies for North America and East Asia.
Regional Model 15 20 25 30 35 40 45
North American w/o noise and Clim. 3.108 3.226 3.345 3.632 3.711 3.843 3.853
w/o noise 3.095 3.144 3.262 3.342 3.406 3.604 3.611
w/o Clim. 2.980 3.027 3.103 3.203 3.250 3.332 3.312
Default 2.715 2.764 2.913 3.007 3.048 3.113 3.123
East Asia w/o noise and Clim. 3.037 3.151 3.275 3.554 3.577 3.751 3.797
w/o noise 2.951 2.943 3.039 3.102 3.153 3.188 3.253
w/o Clim. 2.966 3.099 3.160 3.201 3.332 3.440 3.545
Default 2.591 2.686 2.834 2.919 2.886 2.965 3.011
Additionally, we have supplemented our analysis with ablation experiments across two different regions to further investigate the performance of our model in different places. Based on Table 12, we provide a comprehensive discussion:
  • Due to more complex orography and land–sea contrast, T2m RMSE over North America is overall higher than over East Asia; moreover, in North America, the “w/o Clim.” case (only noise) is better than “w/o noise” (only climatology), indicating that noise perturbations bring larger gains than climatology in more complex terrain.
  • From 15 to 45 days, the advantage of the Default over all ablations persists and generally enlarges, confirming that our components are especially helpful for mitigating long-lead skill degradation while still improving short leads.
Table 13: T2m RMSE of ablation studies for North America and East Asia.
Variables Model 15 20 25 30 35 40 45
T850 w/o noise and Clim. 3.867 3.923 3.883 3.990 4.006 4.062 4.116
w/o noise 3.599 3.614 3.691 3.846 3.875 3.982 3.931
w/o Clim. 3.412 3.495 3.573 3.666 3.666 3.698 3.860
Default 3.249 3.283 3.307 3.346 3.442 3.481 3.563
Wind10 w/o noise and Clim. 2.844 2.861 2.894 2.815 2.919 2.922 3.110
w/o noise 2.694 2.719 2.756 2.793 2.804 2.903 2.923
w/o Clim. 2.658 2.598 2.558 2.670 2.748 2.694 2.795
Default 2.418 2.420 2.422 2.437 2.504 2.532 2.598
From Table 13, for different variables (T850 and Wind10), the Default model (with noise + climatology) consistently achieves the lowest RMSE across all leads, confirming that our design benefits variables with diverse physical scales.
To investigate performance during challenging conditions, we analyzed the model’s results during the severe 2018 North American (NAM) and 2018 East Asian (EA) heatwaves, starting from June 28, 2018, and July 9, 2018, respectively in Table 14.
Table 14: T2m RMSE of ablation studies for North America and East Asia.
Lead Time Model NAM EA
25-day ClimaX 2.948 3.046
FuXi-S2S 2.634 2.851
TianQuan-S2S 2.522 2.721
40-day ClimaX 3.081 3.120
FuXi-S2S 2.724 2.942
TianQuan-S2S 2.685 2.843
We can find that TianQuan-S2S achieves better results compared to existing models (ClimaX, FuXi-S2S) even under these extreme cases, demonstrating its practical usefulness. While these results are encouraging, we acknowledge that improvements on extreme-weather metrics are more modest than for bulk statistics, suggesting that rare, high-impact events in complex terrain remain more challenging and warrant further targeted investigation.
Table 15: T2m RMSE Comparison of Ensemble Mean Results for Different Noise Scale.
Regional Model 15 20 25 30 35 40 45
Input Perturbations $\sigma=0.1$ 2.509 2.554 2.568 2.569 2.624 2.630 2.718
$\sigma=0.5$ 2.477 2.535 2.539 2.543 2.569 2.592 2.679
$\sigma=1$ 2.446 2.506 2.527 2.538 2.556 2.624 2.674
$\sigma=1.5$ 2.551 2.580 2.594 2.614 2.664 2.692 2.734
$\sigma=2$ 2.611 2.637 2.642 2.652 2.696 2.710 2.774
Fixed Noise $\sigma=0.1$ 2.507 2.538 2.548 2.553 2.607 2.614 2.696
$\sigma=0.5$ 2.505 2.556 2.565 2.583 2.617 2.635 2.707
$\sigma=1$ 2.497 2.556 2.560 2.583 2.601 2.616 2.719
$\sigma=1.5$ 2.615 2.638 2.648 2.663 2.695 2.717 2.786
$\sigma=2$ 2.681 2.708 2.721 2.728 2.762 2.798 2.886
Notably, the results shown in Table 15, the performance of Input Perturbation and Fixed Noise rapidly drops when the noise scale increases to 1.5.